An upstretched spring is 10 ft. long. A pull of 40 lb stretches the spring by 1/2 ft. Find the work done in stretching from 10 ft. to 14 ft.


1
Expert's answer
2021-05-03T10:32:47-0400

Spring constant

"K=\frac{40}{\frac{1}{2}}=80lb/ft"


Work done will be store as potential energy of spring so

"W=\frac{1}{2}kx^2"

"W=\frac{1}{2}(80)(4)^2=640 lbft"